# Generic Antiholomorphic Polymonial Vector Fields

### v1

@@ Abstract.
An analytic classification of generic anti-polynomial vector fields $\dot{z}=\overline{P(z)}$ is given in term of a topological and an analytic invariant. The number of generic strata in the parameter space is counted for each degree of $P$. A Realization Theorem is established for each pair of topological and analytic invariants. The non-generic case of a maximal number of heteroclinic connections is also given a classification. The bifurcation diagram for the quadratic case is presented.
@@ 1. Introduction
In this paper, we are mainly concerned with the study and classification of vector fields $\overline{P(z)}{\partial\partial z}$, where $P$ is a polynomial (and $\overline{\,\cdot\,}$ is the complex conjugation). By classifying, we mean describing the equivalence classes of the family of vector fields $\overline{P(z)}{\partial\partial z}$ under some equivalence relation, such as the topological orbital equivalence or a holomorphic or antiholomorphic equivalence.
There are several motivations for this study. Early in the development of complex analysis, complex potentials found many applications in physics; broadly speaking, given a (possibly ramified) holomorphic function $f\colon U\to\mathbb{C}$ (the complex potential), over some open set $U$, the dynamics of $\overline{f^{\prime}(z)}{\partial\partial z}$ may be thought of as the laminar flow of a Newtonian fluid in the plane. Although the flow of a real fluid is much more complicated than this, these simple potential flows can be used to construct a variety of flows, by superposition, patching or otherwise, that are much closer to what is observed in the physical world. Potential flows have been used in the early development of aerodynamics [VD75] and in recent work (e.g. [CKM+17]). From this point of view, (generic) anti-holomorphic polynomials arise as flows on $\mathbb{C}$ for which the complex potential
  • is meromorphic on the whole Riemann;
  • and has only simple zeroes in $\mathbb{C}$ and no homoclinic loops or heteroclinic connections.
Such a flow with $k+1$ stagnation point is characterized at the topological level by the organization of $k$ strips (zone isomorphic to horizontal strip) in which the fluid flows from infinity to infinity. In Section 3, we prove that there $A^{r}(k+2)$ such configurations, where $A^{r}(n)$ is the number of non crossing trees with $n$ vertices modulo rotations of order $n$. Then the circulation and the flux of each strip determines analytically the polynomial.
There has also been recent development in approximation of potential flows in planar domains [Bad20]. Rational functions are used to approximate the flow. The method uses a cluster of poles near the corner to achieve very fast approximation when the number of poles increases exponentially. Recent work [Tre24] on the approximation of solutions of the Laplace equation $\Delta f=0$ in planar domains showed that rational approximations performs better than polynomial approximations on some non-convex regions.
A structurally stable anti-polynomial (i.e. a polynomial in ${\overline{z}}$) vector field has invariant regions isomorphic to a horizontal strip. The circulation and the flux across a transversal section of this region is an analytic invariant that helps to classify those vector fields.
@@
On a different note, we also see this work as part of the analysis of (holomorphic) polynomial vector fields on $\mathbb{C}$. Indeed, it is well known that such a vector field has a pole at infinity, which is the organizing centre of the dynamics. Those vector fields have been well-studied in [DES05] and then by [BD10]. To study the unfoldings of this singularity, it is natural to consider ${1P(z)}{\partial\partial z}$, which leads to $\overline{P(z)}{\partial\partial z}$ by observing that
@@
The paper is organized as follow. In Section 2, we study the separatrix graph and the two types of components in the complement of this graph. In Section 3, we discuss the topological properties of the phase portrait and we count the number of equivalence classes under topological orbital equivalence. Section 4 is devoted to the analytic invariants of the vector fields. In Section 5, we prove a realization theorem, proving that every phase portrait is completely determined by the topological and analytic invariants. Lastly, in Section 6, we initiate the study of non-structurally stable vector fields.
@@ 2. Description of the phase portrait
We consider the system
$\dot{z}=\overline{P(z)}$ (1)
where $P$ is a monic and centred polynomial of degree $k+1$. Near infinity, the behaviour of the dynamics is of the form $\dot{w}=-\frac{w^{k+3}}{|w|^{2(k+1)}}(1+o(1))$, with $w=\frac{1}{z}$ and $|w|$ small (see Figure 1). Hence, the system $\dot{w}$ has a parabolic point of codimension $k+2$ near the origin.
Through the paper, we will denote by $\Xi_{k+1}$ the set of such vector fields and we will implicitly identify any such vector field with its defining polynomial.
Figure 1. Phase portrait of $\dot{z}=\overline{P(z)}$ outside a large enough disk, with $\deg P=5$
The phase portrait may contain a finite number of simple saddle points, multiple saddle points and heteroclinic connections. The latter two are the only possible non-structurally stable objects. Indeed, any singular point must be a saddle point, so no nodes, parabolic points or centres are possible, and no homoclinic loops or limit cycles can exist, since they must have a positive index number. For the same reason, a domain cannot be bounded by orbits.
The basin of attraction of $\infty$ consists of a finite number of connected and simply connected regions of $\mathbb{C}\setminus E(\Gamma_{P})$ bounded by incoming orbits and singular points. If we know the basin of attraction at $\infty$ and its boundary, we can reconstruct the whole phase portrait. This allows us to reduce our problem to the study of the separatrix graph (defined in the following paragraph).
@@ 2.1. Separatrix graph
The separatrix graph, formed of singular points and separatrices, was already used in [DES05] in the case of holomorphic polynomial vector fields. So it is no surprise that this object completely determines the topological structure of the phase portrait in the case at hand. At the end of Section 2, we will reduce the study of the structurally stable phase portrait under topological orbital equivalence to that of different graphs, deduced from the separatrix graph, called noncrossing trees. The discrete and combinatorial properties of those trees were already studied in [Noy98].
@@
    By abuse of language, we will use the term separatrix both for incoming and outgoing separatrices.
@@
    We will call separatrix graph of (1) the graph $\Gamma_{P}$ on the zeros of $P$ with the separatrices as edges.
@@
Here is the motivation for this terminology. A petal zone contains a petal of the parabolic point at infinity and is bounded by a single curve formed by the union of separatrices and single points. A sepal zone is a domain isomorphic to a strip that fits in the interstice spaces between petals. Note that in the rectifying coordinate $t=\int P$, a sepal zone becomes a horizontal strip, and petal zone, an upper or lower half-plane.
@@
Then the orbit of $\gamma(\tau_{0})$ must land on the boundary so that the orbit of $t^{-1}(\gamma(\tau_{0}))$ must land on one the boundaries, a contradiction.
@@ 2.2. Structurally stable vector fields
We now specialize our study to the case of structurally stable vector fields, from here until Section 6. Broadly speaking, a vector field is structurally stable if its phase portrait remains topologically the same after a small perturbation. In the case of antiholomorphic polynomial vector fields, this is equivalent to having only simple singular points and no heteroclinic connection. (Limit cycles and homoclinic loops were already impossible in these systems.)
Now, consider the compactification ¯$\overline{\mathbb{C}}=\mathbb{C}\cup S^{1}$ of the complex plane given by a gnomonic projection on a half-sphere, so that $\mathbb{C}$ is mapped on the lower half-sphere and the circle at infinity is mapped on the equator. (Note that this projection is not conformal, however we only use it for topological consideration.) We can divide the circle in $k+2$ attracting intervals, where orbits are landing at infinity, alternating with $k+2$ repulsive intervals, where orbits are leaving infinity. For convenience, we will identify each interval to a point $\exp\!\left({2ij\pi 2k+4}\right)$, with an attractive (resp. repulsive) interval for $j$ even (resp. $j$ odd).
@@
At each zero $z_{j}$ of $P$, there are $2$ incoming (solid on Figure 3) and $2$ outgoing (dashed on Figure 3) orbits. These come in alternating order since every singularity is a saddle point. By the previous discussion, each incoming (resp. outgoing) separatrix must come from one of the $k+2$ directions escaping (resp. going toward) the point $\infty$. Taking orientation into consideration, this gives $42k+$ 4 points on $S^{1}$ and, identifying opposite separatrices to a single curve, $22k+$ 2 edges (see Figure 3).
Figure 3. The phase portrait of (1) in a disk $B_{R}(0)$ for $R>>1$ with $k=2$.
This data organizes the topological behaviour of the phase portrait. In fact, knowing only half of this information is sufficient. Indeed, if we are only given the information from the incoming separatrices, we can completely recover the (topological) data of the outgoing ones since singular points are saddle points and incoming separatrices determine the basin of attraction at $\infty$ (see Figure 3). In a similar fashion, the outgoing separatrices also determine the phase portrait.
We encode this data in a graph $G$ embedded in $\overline{\mathbb{C}}$, having $k+2$ vertices on $\partial\overline{\mathbb{C}}$ and $k+1$ edges. Note that if we consider both the incoming and outgoing graphs, we get the separatrix graph of the system (1). Remark also that given an incoming (resp. outgoing) graph, its associated outgoing (resp. incoming) graph is (isomorphic to) its dual (for the definition of the dual graph, we refer the reader to Definition 3.5). Hence, we’ll refer to the incoming graph as the combinatorial invariant of (1). This choice is arbitrary, i.e. it would have been equivalent to give the outgoing graph as the combinatorial invariant. We resume this discussion in the following results.
@@
The dual $G^{*}$ of an incoming graph $G$ is its outgoing graph and $G^{**}=G$. Moreover, the separatrix graph is given by $G$ and $G^{*}$.
@@
If $G$ is an incoming graph, then it’s embedded (i.e. no two vertices intersect).
@@
If $G$ is an incoming graph, then $G$ has no cycle.
@@
By contradiction, assume that there exists a cycle $c$ and take any edge $e$ in the cycle. The edge $e$ corresponds to the two separatrices entering some zero $z_{j}$. Now recall that there must exists an outgoing separatrix $\gamma$ at $z_{j}$ that satisfies $\lim_{t\to\infty}|\gamma(t)|=\infty$. But any point $\gamma(t_{0})$ is in the basin of attraction of one of the $z_{j}$’s, giving a contradiction (see Figure 4).
@@
An incoming graph $G$ has $k+1$ edges, so it divides $\overline{\mathbb{C}}$ in $k+2$ connected, simply-connected components. In each component, there is a vertex of the dual tree $G^{*}$ on the circle at infinity. If $n_{j}$ is the incidence of the $j$-th vertex of the dual, it means this connected component gets divided into $n_{j}+1$ connected components by the edges of the dual tree. Since the dual is the outgoing tree, it also has $k+1$ edges, so that $\sum n_{j}=2(k+1)$. Therefore, the number of connected components of $\mathbb{C}\setminus\Gamma_{P}$ is $\sum(n_{j}+1)=2(k+1)+k+2=3k+4$.
Moreover, for each $n_{j}$, the corresponding connected component is divided in two petal zones and $n_{j}-1$ sepal zones. Therefore, there are $\sum_{j}(n_{j}-1)=k$ sepal zones and $42k+$ 4 petal zones. ∎
@@ 3. Combinatorial invariant
A noncrossing tree of order $n$ (or $n$-nc tree for short) is a tree on $n$ vertices and $n-1$ edges that is embedded in $\overline{\mathbb{D}}$ such that every vertex belongs in $\partial\mathbb{D}$ and no two embedded vertices cross.
Denote by $\sim_{\text{top}}$ the relation of topological orbital equivalence on $\Xi_{k+1}$. By Lemmas 2.5 and 2.6, counting incoming graphs amounts to counting some nc trees. We have the following.
@@
Let $k\geq 2$. There is a bijection between $(k+2)$-nc trees and the equivalence classes of $\Xi_{k+1}/\sim_{\text{\rm top}}$.
Proof.
One direction of the proof of Theorem 3.2 is given by the two lemmas that characterize the phase portrait of antiholomorphic polynomial vector fields. The other direction is a corollary of the Realization Theorem 5.1 which says that for each nc tree, there is a vector field having this tree as a combinatorial invariant. The latter theorem will be discussed in Section 5. ∎

### v2

@@ Abstract.
An analytic classification of generic anti-polynomial vector fields $\dot{z}=\overline{P(z)}$ is given in term of a topological and an analytic invariants. The number of generic strata in the parameter space is counted for each degree of $P$. A Realization Theorem is established for each pair of topological and analytic invariants. The non-generic case of a maximal number of heteroclinic connections is also given a classification. The bifurcation diagram for the quadratic case is presented.
@@ 1. Introduction
In this paper, we are mainly concerned with the study and classification of vector fields $\overline{P(z)}{\partial\partial z}$, where $P$ is a polynomial and $\overline{\,\cdot\,}$ is the complex conjugation. By classifying, we mean describing the equivalence classes of the family of vector fields $\overline{P(z)}{\partial\partial z}$ under some equivalence relation, such as the topological orbital equivalence or holomorphic equivalence.
Early in the development of complex analysis, anti-holomorphic vector fields found many applications in physics because of complex potentials; broadly speaking, given a (possibly ramified) holomorphic function $f\colon U\to\mathbb{C}$ (the complex potential), over some open set $U\subseteq\mathbb{C}$, the dynamics of $\overline{f^{\prime}(z)}{\partial\partial z}$ may be thought of as the laminar flow of a Newtonian fluid in the plane. Although the flow of a real fluid is much more complicated than this, these simple potential flows can be used to construct a variety of flows, by superposition, patching or otherwise, that are much closer to what is observed in the physical world. Potential flows have been used in the early development of aerodynamics [VD75] and in recent work (e.g. [CKM+17]). From this point of view, generic anti-holomorphic polynomials arise as vector fields on $\mathbb{C}$ that are
  • meromorphic on the whole Riemann sphere;
  • and have only simple singularities (forced to be saddle points) in $\mathbb{C}$ and no homoclinic loops or heteroclinic connections.
Such a flow with $k+1$ saddle points is characterized at the topological level by the organization of $k$ strips (zone isomorphic to horizontal strip) in which the fluid flows from infinity to infinity. In Section 3, we prove that there are $A^{r}(k+2)$ such configurations (see Equation (8)), where $A^{r}(n)$ is the number of non crossing trees with $n$ vertices modulo rotations of order $n$. Then the circulation and the flux of each strip determines completely the polynomial.
There has also been recent development in approximation of potential flows in planar domains [Bad20]. Rational functions are used to approximate the flow. The method uses a cluster of poles near the corner to achieve very fast approximation when the number of poles increases exponentially. Recent work [Tre24] on the approximation of solutions of the Laplace equation $\Delta f=0$ in planar domains showed that rational approximations perform better than polynomial approximations on some non-convex regions.
@@
On a different note, we also see this work as part of the analysis of (holomorphic) polynomial vector fields on $\mathbb{C}$. Indeed, it is well known that such a vector field has a pole at infinity, which is the organizing centre of the dynamics. To study the unfoldings of this singularity, it is natural to consider ${1P(z)}{\partial\partial z}$, which leads to $\overline{P(z)}{\partial\partial z}$ by observing that
@@
The paper is organized as follow. In Section 2, we study the separatrix graph and the two types of components in the complement of this graph. We also discuss the topological properties of the phase portrait. In Section 3, we count the number of equivalence classes under topological orbital equivalence. Section 4 is devoted to the analytic invariants of the vector fields. In Section 5, we prove a realization theorem, proving that every phase portrait is completely determined by the topological and analytic invariants. Lastly, in Section 6, we initiate the study of non-structurally stable vector fields.
@@ 2. Description of the phase portrait
A polynomial $P(z)=a_{k+1}z^{k+1}+\cdots+a_{0}$ is monic if $a_{k+1}=1$ and centred if $a_{k}=0$.
Notation 2.1.
Through the paper, we will denote by $\Xi_{k+1}$ the set of monic, centred anti-polynomial vector fields of degree $k+1$ and we will implicitly identify any such vector field with its defining polynomial. We will note such a vector field $\overline{P}{\partial\partial z}$, with
$\overline{P(z)}=\overline{z^{k+1}+a_{k-1}z^{k-1}+\cdots+a_{0}}.$
For $\overline{P}\in\Xi_{k+1}$, we consider the system
$\dot{z}=\overline{P(z)}.$ (1)
Every singular point must a saddle point; this can be checked directly in polar coordinates in a neighbourhood of the singular point. Therefore, no nodes, parabolic points or centres are possible in $\mathbb{C}$. Homoclinic loops or limit cycles cannot occur, since they must have a positive index number. For the same reason, a domain cannot be bounded by orbits.
Near infinity, the behaviour of the dynamics is of the form $\dot{w}=-\frac{w^{k+3}}{|w|^{2(k+1)}}(1+o(1))$, with $w=\frac{1}{z}$ and $|w|$ small. Hence, the system $\dot{w}$ has a parabolic point of codimension $k+2$ near the origin. Consider the blow-up ¯$\overline{\mathbb{C}}=\mathbb{C}\sqcup S^{1}$ at $\infty$ of the Riemann sphere $\hat{\mathbb{C}}$ with fiber $S^{1}$ (one point for every oriented direction at infinity). In this space, we can make sense of the notion of going to $\infty$ in the direction $\theta$ for every $\theta\in S^{1}$. The blown up parabolic point at infinity has $k+2$ attracting points and $k+2$ repulsing points on $S^{1}$ in alternating order. These marked points are located at $\exp\!\left({2ij\pi 2k+4}\right)$ on the circle at infinity, with an attractive (resp. repulsive) point for $j$ even (resp. $j$ odd). See Figure 1.
(a) Outside a large disk in $\mathbb{C}$
(b) Near the circle at infinity in $\overline{\mathbb{C}}$
Figure 1. Phase portrait of $\dot{z}=\overline{P(z)}$, with $\deg P=5$
Definition 2.2.
Two vector fields $\overline{P_{1}}{\partial\partial z}$ and $\overline{P_{2}}{\partial\partial z}$ are said to be topologically orbitally equivalent if there exists a homeomorphism $h:\mathbb{C}\to\mathbb{C}$ that maps orbits of $\overline{P_{1}}{\partial\partial z}$ into orbits of $\overline{P_{2}}{\partial\partial z}$. In this paper, we further require that the homeomorphism fixes the order of the marked points at infinity (see above).
The phase portrait may contain a finite number of simple saddle points, multiple saddle points and heteroclinic connections. The latter two are the only possible non-structurally stable objects.
@@ 2.1. Separatrix graph
The separatrix graph, formed of singular points and separatrices, was already used in [DES05] in the case of holomorphic polynomial vector fields, so it is no surprise that this object completely determines the topological structure of the phase portrait in the case at hand. At the end of Section 2, we will reduce the study of the structurally stable phase portrait under topological orbital equivalence to that of different graphs, deduced from the separatrix graph, called noncrossing trees. The discrete and combinatorial properties of those trees were already studied in [Noy98].
The basin of attraction of $\infty$ consists of a finite number of connected and simply connected regions of $\mathbb{C}$ bounded by singular points and orbits with bounded $\omega$-limits. If we know the basin of attraction at $\infty$ and its boundary, we can reconstruct the whole phase portrait. This allows us to reduce our problem to the study of the separatrix graph (defined in the following paragraph).
@@
    An incoming (resp. outgoing) separatrix is an orbit of a non singular point on the boundary of the basin of attraction (resp. repulsion) of the point at infinity.
@@
    We will call separatrix graph of $\overline{P}\in\Xi_{k+1}$ the graph $\Gamma_{P}$ on the zeros of $P$ with the separatrices as edges.
@@
Here is the motivation for this terminology. A petal zone contains a petal of the parabolic point at infinity and is bounded by a single curve formed by the union of separatrices and singular points. A sepal zone is a domain isomorphic to a strip that fits in the interstice spaces between petals. Note that in the rectifying coordinate $t=\int P$, a sepal zone becomes a horizontal strip, and petal zone, an upper or lower half-plane.
@@
Then the orbit of $\gamma(\tau_{0})$ must land on the boundary so that the orbit of $t^{-1}(\gamma(\tau_{0}))$ must land on one the boundaries, a contradiction. ∎
@@ 2.2. Structurally stable vector fields
We now specialize our study to the case of structurally stable vector fields, from here until Section 6.
Broadly speaking, a vector field is structurally stable if its phase portrait remains topologically the same after a small perturbation. In the case of antiholomorphic polynomial vector fields, this is equivalent to having only simple singular points and no heteroclinic connection. (Limit cycles and homoclinic loops were already impossible in these systems.)
Notation 2.6.
We note $\Xi_{k+1}^{\prime}$ the subspace of generic monic centred anti-polynomial vector fields in $\Xi_{k+1}$.
@@
At each zero $z_{j}$ of $P$, there are $2$ incoming and $2$ outgoing separatrices (see Figure 3(a)). These come in alternating order since every singularity is a saddle point.
Given a zero $z_{j}$ of $P$, define an edge as the union of $z_{j}$ and its two incoming separatrices. Now, embedding these edges in $\overline{\mathbb{C}}$ taking the endpoints as vertices, this gives a graph embedded in $\overline{\mathbb{C}}$, with $k+2$ vertices and $k+1$ noncrossing edges (see Figure 3(b)). Call this graph the incoming graph of $P$. This data organizes the topological behaviour of the phase portrait since singular points are saddle points and incoming separatrices determine the basin of attraction at $\infty$ (see Figure 3). This will be proved in Theorem 3.10.
(a) Separatrices (incoming in solid, outgoing in dashed) and singular points in $\overline{\mathbb{C}}$
(b) Incoming graph
Figure 3. Incoming graph obtained from a separatrix graph for $k=2$.
We’ll refer to this graph as the combinatorial invariant of $\overline{P}{\partial\partial z}$. It would have been equivalent to consider edges as union of zeroes and outgoing separatrices instead, which we call the outgoing graph, as the combinatorial invariant. Indeed, we have the following result.
@@
The dual of an incoming graph $G$ is isomorphic to the outgoing graph. Moreover, the separatrix graph is given by the union of the incoming graph and the outgoing graph in $\overline{\mathbb{C}}$.
Proof.
Recall that $\mathbb{C}\setminus G$ is the basin of attraction of infinity. If two components share a boundary, then there is singular point in this boundary and an outgoing separatrix going in each component, forming an edge of the outgoing graph. If $G^{\prime}$ denote the outgoing graph, we can do a symmetric argument on $\mathbb{C}\setminus G^{\prime}$. ∎
@@
If $G$ is an incoming graph, then it is embedded (i.e. no two vertices intersect).
@@
If $G$ is an incoming graph, then $G$ has no cycles.
@@
By contradiction, assume that there exists a cycle $c$ and take any edge $e$ in the cycle. The edge $e$ corresponds to the two separatrices entering some zero $z_{j}$. Let $U\subseteq\overline{\mathbb{C}}$ be the region bounded by the edges of the cycle $c$. Recall that there must exist an outgoing separatrix $\gamma$ at $z_{j}$ that satisfies $\lim_{t\to\infty}|\gamma(t)|=\infty$. But the only ends of the region $U$ correspond to directions going away from infinity, a contradiction (see Figure 4).
@@
An incoming graph $G$ has $k+1$ edges, so it divides $\overline{\mathbb{C}}$ in $k+2$ connected, simply-connected components. In each component, there is a vertex of the dual tree $G^{*}$ on the circle at infinity. If $n_{j}$ is the incidence of the $j$-th vertex of the dual, it means this connected component gets divided into $n_{j}+1$ connected components by the edges of the dual tree. Since the dual is the outgoing tree, it also has $k+1$ edges, so that $\sum_{j=1}^{k+2}n_{j}=2(k+1)$. Therefore, the number of connected components of $\mathbb{C}\setminus\Gamma_{P}$ is $\sum_{j}(n_{j}+1)=2(k+1)+k+2=3k+4$.
Moreover, for each $n_{j}$, the corresponding connected component is divided in two petal zones and $n_{j}-1$ sepal zones. It follows that there are $\sum_{j}(n_{j}-1)=k$ sepal zones and $42k+$ 4 petal zones. ∎
@@ 3. Combinatorial invariant
By Lemmas 2.8 and 2.9, the incoming and outgoing graphs are part of the family of graphs known as noncrossing trees in the literature.
@@
A noncrossing tree of order $n$ (or $n$-nc tree for short) is a rooted tree on $n$ vertices and $n-1$ edges that is embedded in $\overline{\mathbb{D}}$ ($\cong\overline{\mathbb{C}}$), where $\mathbb{D}$ is the unit disk in $\mathbb{C}$, such that every vertex belongs in $\partial\mathbb{D}$ and no two embedded vertices cross.
We ask that an isomorphism of $n$-noncrossing trees fixes the root and preserves the cyclic order of the vertices in the boundary circle.
Denote by $\sim_{\text{top}}$ the relation of topological orbital equivalence fixing the marked point at infinity (see Definition 2.2) on $\Xi_{k+1}$. Recall that $\Xi_{k+1}^{\prime}$ is the subspace of monic centred generic anti-polynomial vector fields.
@@
For each $k\geq 0$, there is a bijection between $(k+2)$-nc trees and the equivalence classes of $\Xi_{k+1}^{\prime}/{\sim_{\text{\rm top}}}$.
It will be a corollary of Theorem 5.1 of Section 5 that every nc tree can be realized as the combinatorial invariant of some generic antiholomorphic polynomial vector field. Hence, counting the number of orbital topological equivalence classes amounts to counting the number of noncrossing trees.